OpenStudy (anonymous):

How to do you add and subtract exponents? Ex. 2x squared minus x squared

OpenStudy (anonymous):

2x^2 - 1x^2 = x^2

OpenStudy (anonymous):

treat it the same way you would x

OpenStudy (anonymous):

Read up on your exponent laws. Like items can be added or subtracted like normal terms. (2x)^2 - (x)^2 = (x)^2